OpenStudy (anonymous):

5w+6=9-w help full instructions soooo confused

jhonyy9 (jhonyy9):

6w=3 w=3/6 w/1/2

jhonyy9 (jhonyy9):

w=1/2

jhonyy9 (jhonyy9):

so

jhonyy9 (jhonyy9):

sorry

OpenStudy (anonymous):

loool its cool but can you explain it to me in detail please ?

OpenStudy (anonymous):

5w+6=9-w send w to the left and 6 to the right side 5w+w=9-6 6w=3 w=1/2

OpenStudy (anonymous):

still confused lool making progress though tahanks